OpenStudy (anonymous):

Field goals worth three points and touchdowns worth seven points were scored during the homecoming football game. With 11 scores, the team set a county record of 57 points. Build a system of equations that can be used to find how many field goals and touchdowns were made.

OpenStudy (wolf1728):

fg + td = 11 3 * fg + 7 * td = 57 So 5 field goals (5 * 3 = 15) + 6 touchdowns (6 * 7 = 42)